At the end of its third year of operations, the Sandifer manufacturing Co had $ 4,510,000 in revenue, $ 3,334,000 in cost of good sold $ 447,000 in operating expenses which included depreciation expense of $ 140,000 and a tax liability equal to 34% of the firm taxable income. What is the net income of the firm for the year.
Answer #1

Net Income = (Revenue - COGS - Operating Expenses - Depreciation)*(1-Tax Rate) = (4,510,000 - 3,334,000 - 447,000 - 140,000)*(1-34%) = 338,740.00

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well